# EDBench: Large-Scale Electron Density Data for Molecular Modeling

### v1

@@ Abstract
Existing molecular machine learning force fields (MLFFs) generally focus on the learning of atoms, molecules, and simple quantum chemical properties (such as energy and force), but ignore the importance of electron density (ED) $\rho(r)$ in accurately understanding molecular force fields (MFFs). ED describes the probability of finding electrons at specific locations around atoms or molecules, which uniquely determines all ground state properties (such as energy, molecular structure, etc.) of interactive multi-particle systems according to the Hohenberg-Kohn theorem. However, the calculation of ED relies on the time-consuming first-principles density functional theory (DFT) which leads to the lack of large-scale ED data and limits its application in MLFFs. In this paper, we introduce EDBench, a large-scale, high-quality dataset of ED designed to advance learning-based research at the electronic scale. Built upon the PCQM4Mv2, EDBench provides accurate ED data, covering 3.3 million molecules. To comprehensively evaluate the ability of models to understand and utilize electronic information, we design a suite of ED-centric benchmark tasks spanning prediction, retrieval, and generation. Our evaluation on several state-of-the-art methods demonstrates that learning from EDBench is not only feasible but also achieves high accuracy. Moreover, we show that learning-based method can efficiently calculate ED with comparable precision while significantly reducing the computational cost relative to traditional DFT calculations. All data and benchmarks from EDBench will be freely available, laying a robust foundation for ED-driven drug discovery and materials science.
@@ 1 Introduction
With the widespread adoption of deep learning in molecular dynamics (MD) simulations, machine learning force fields (MLFFs) have become efficient and promising computational tools, significantly advancing research in physics, chemistry, biology, and materials science (Wu et al., 2023; Kabylda et al., 2023; Frank et al., 2024). State-of-the-art MLFFs methods typically employ geometric deep learning to model atomic interactions within molecules, a strategy that has proven to be effective (Zheng et al., 2025). These models are generally built upon many-body interactions at the atomic level, including one-body (atomic attributes such as types and coordinates (Satorras et al., 2021)), two-body (interatomic distances (Schütt et al., 2018; Kabylda et al., 2023)), three-body (bond angles (Gasteiger et al., 2020; Schütt et al., 2021)), four-body (torsions (Gasteiger et al., 2021; Liu et al., 2022; Wang et al., 2024) and improper torsions (Zheng et al., 2025)), and five-body interactions (Wang et al., 2023b).
@@
Although existing MLFFs have demonstrated great potential in modeling molecular force fields (MFFs), they primarily focus on capturing coarse-grained, atom-level many-body interactions (Liu et al., 2023; Zheng et al., 2025), while often overlooking the critical role of microscopic electron distribution in understanding molecular interactions (Schuch and Verstraete, 2009; Gao et al., 2019; Pederson et al., 2022). It is well known that the spatial distribution of electrons directly influences the interactions between atoms within a molecule, providing the most direct and fundamental information for interpreting molecular force fields (Hapala et al., 2016). Electron density (ED), as a fundamental physical quantity in quantum mechanics that describes the spatial distribution of electrons, offers a more fine-grained and physically grounded representation of molecular systems according to Hohenberg–Kohn (HK) theorem (Hohenberg and Kohn, 1964). Therefore, explicitly incorporating ED into the modeling process holds promise for bridging the gap between microscopic electronic behavior and macroscopic force fields, further improving both the accuracy and generalization of MLFFs. Therefore, as illustrated in Figure 1(a), the primary objective of this work is to advance current MLFFs beyond the atom-level learning paradigm toward electron-level modeling, enabling a more accurate and physically grounded description of molecular interactions.
@@
However, advancing MLFFs toward an electron-level understanding faces two major challenges: (i) the lack of large-scale, high-quality ED datasets, which are essential for pretraining and could fundamentally reshape the paradigm of MLFFs modeling. (ii) the absence of an ED-centric benchmark to systematically explore the feasibility and effectiveness of ED-based modeling frameworks. Specifically, the acquisition of ED data can be categorized into two approaches: experimental methods (such as X-ray diffraction (Casati et al., 2016; Kasai et al., 2018), electron diffraction (Gruene et al., 2021)) and theoretical calculation methods. Due to the reliance on expensive physical equipment, experimental methods inevitably limit data acquisition, making theoretical methods more popular. Theoretical calculations typically use density functional theory (DFT) (Bartolotti and Flurchick, 1996; Medvedev et al., 2017), the most common approach, to compute the ED of molecules. Although DFT does not depend on specialized observation equipment, its calculations are highly computationally intensive and time-consuming, making the acquisition of large-scale, high-quality ED datasets particularly difficult. In addition, the MLFFs community is still in the early stages of learning effective representations from ED, which makes the development of an ED-based evaluation protocol particularly important for the rapid advancement of ED representation learning.
@@
To address the two key challenges outlined above, we introduce EDBench, a large-scale and high-fidelity dataset of ED, as shown in Figure 1(b). Following Jacob’s ladder (Sousa et al., 2007), as shown in Figure 1(c), we adopt higher-rung hybrid functionals as the underlying DFT methods to ensure the quality of the EDBench dataset. Specifically, we use B3LYP function combined with 6-31G**/+G ** basis set and run over 205,000 core-hours on a high-performance computer to generate EDBench. EDBench comprises 3,359,472 drug-like molecules with corresponding ED distributions and a comprehensive set of quantum chemical properties, including energy components, orbital energies, and multipole moments, thus providing a solid foundation for systematically investigating the role of ED in molecular modeling. To enable rigorous benchmarking and model development, we further design an ED-centric benchmark suite covering three task categories: (i) quantum property prediction, including four core tasks—energy components prediction (ED5-EC), orbital energe estimation (ED5-OE), multipole moment regression (ED5-MM), and open-/closed-shell classification (ED5-OCS)—to evaluate how well ED alone can serve as a sufficient descriptor for inferring fundamental quantum properties; (ii) cross-modal retrieval between molecular structures and ED (ED5-MER), designed to probe the mutual consistency and representational alignment between structural and density spaces, which is critical for density-based force field construction and virtual screening; and (iii) ED prediction from molecular structures (ED5-EDP), aimed at approximating DFT-level density fields at significantly reduced computational cost, thereby enabling scalable quantum-aware modeling. Finally, we evaluate several state-of-the-art deep learning models on the proposed benchmark, offering the first large-scale assessment of ED understanding in data-driven systems.
@@ 2.1 Density Functional Theory (DFT)
The quantum mechanical description of many-electron systems is one of the core issues in modern physics and chemistry. Schrödinger equation (Schrödinger and E., 1926) as the fundamental equation of quantum mechanics, is challenging to solve directly. Consequently, researchers introduced various wave function-based approximation methods to simplify the problem, such as, Born–Oppenheimer (Combes et al., 1981) and Hartree-Fock method (Szabo and Ostlund, 1996). Those methods scale with the number of electrons $n$ as $\mathcal{O}(n^{4})$ or more, its computational cost remains prohibitive for large polyatomic molecules (More details about the development of quantum mechanics see Appendix A.1.1). In contrast, Density Functional Theory (DFT) is more suitable for complex systems due to its lower computational cost ($\mathcal{O}(n^{3})$) and incorporation of electron correlation effects (Gaitan and Nori, 2009). The core concept of DFT is to use electron density (ED) as the fundamental variable instead of the wave function. The Hohenberg-Kohn theorem is the cornerstone of DFT (Hohenberg and Kohn, 1964), which states that the external potential field and the ground-state energy can be completely determined by ED. Thus, by solving for the ED distribution $\rho(r)$ that achieves the lowest energy, the properties of the stable system can be confirmed. The ED $\rho(r)$ can be expressed as:
@@
where $\rho_{\alpha}(\mathbf{r})$ and $\rho_{\beta}(\mathbf{r})$ are the density of $\alpha$-spin electrons and $\beta$-spin electrons. This concept is concretely realized in the Kohn-Sham equations, which transforms the polyelectron system with interactions into single-electron system without interaction, and adds interactions among electrons to exchange-correlation potential (Kohn and Sham, 1965). The Kohn-Sham equations is shown as:
@@
The solution of the Kohn-Sham equations is typically achieved through self-consistent field (SCF) iterations, as shown in the figure 1(b). Initially, a set of initial electron densities $\rho(\mathbf{r})$ is selected, and the effective potential $V_{\text{eff}}(\mathbf{r})$ is calculated based on this initial guess. The Kohn-Sham equations are then solved to obtain new single-electron orbital wave functions $\psi_{i}(\mathbf{r})$ and energies $\varepsilon_{i}$, which are used to update the ED $\rho(\mathbf{r})$. This process is repeated until convergence is achieved, yielding the ED $\rho(\mathbf{r})$ and simultaneously stabilizing the total energy $E$. The choice of the exchange-correlation functional determines the accuracy of DFT, as shown in Figure 1(c). Common approximations include the Local Density Approximation (LDA) (Perdew et al., 1996), the Generalized Gradient Approximation (GGA) (Becke, 1988), and hybrid functionals (B3LYP) (Lee et al., 1988). In this paper, we used B3LYP functional combined with 6-31G**/+G** basis set, which achieves a balance between precision and efficiency(Details on basis set and functional see Appendix A.1.3).
@@ 2.2 Molecular Geometry Learning in Quantum Chemistry
Geometric Deep Learning (GDL) has become a dominant approach for modeling machine learning force fields (MLFFs), primarily focusing on atom-level information such as atomic attributes and interatomic interactions. Specifically, GDL models are built upon first-order atomic features, including atom types and 3D coordinates (Satorras et al., 2021; Cui et al., 2024). To capture geometric relationships while preserving physical consistency, GDL methods incorporate symmetries such as rotational and translational invariance in 3D space (Musil et al., 2021; Gebauer et al., 2022). Consequently, a wide range of models have been developed with built-in invariance or equivariance to Euclidean group E(3) (Satorras et al., 2021) or special Euclidean group SE(3) (Ganea et al., 2021; Schneuing et al., 2024), ensuring that predictions are physically meaningful. Given that atomic interactions—such as bonding, repulsion, and van der Waals forces—play a crucial role in molecular fields, modern GDL methods further incorporate second-order geometric features, including interatomic distances (Unke and Meuwly, 2019; Fuchs et al., 2020), bond types (Fang et al., 2022), and spatial neighborhood structures (Zhang and Bai, 2024). To more precisely capture local structural features, some approaches even extend to higher-order geometric relations such as bond angles (three-body interactions) (Gasteiger et al., 2020; Schütt et al., 2021) and torsional angles (four-body interactions) (Gasteiger et al., 2021; Liu et al., 2022; Wang et al., 2024), thereby improving the expressiveness and accuracy of force field modeling. In contrast to prior works that focus primarily on atom-level representations, our proposed EDBench introduces a large-scale dataset of electronic density (ED), laying the foundation for extending molecular modeling from the atomic scale to the electronic scale. It also provides a new platform and evaluation benchmark for developing GDL methods tailored to electronic structure modeling.
@@ 2.3 Quantum Chemistry (QC) Datasets
To more efficiently predict quantum chemical properties and enhance applications using machine learning models, numerous QC datasets have been constructed to provide rich and reliable data support, as shown in Table 1. QM7 (Rupp et al., 2012) and QM9 (Ramakrishnan et al., 2014) are classic QC datasets, built based on GDB-13 (Blum and Reymond, 2009) and GDB-17 (Ruddigkeit et al., 2012), respectively. QM7-X (Hoja et al., 2021) covers up to 42 physicochemical properties. PubQChemQC (Nakata and Maeda, 2023) offers 85 million ground-state molecular structures and HOMO-LUMO gaps. Additionally, MD17 (Chmiela et al., 2017), MD22 (Chmiela et al., 2023) and WS22 (Jr. et al., 2023) offer force data from molecular dynamics trajectories. Based on QM9 dataset, QH9 (Yu et al., 2023) focuses on the Hamiltonian matrix, and MultiXC-QM9 (Nandi et al., 2023) includes information on chemical reactions. Compared with QC datasets that provide energy (Jie et al., 2019) and force, the proportion of existing ED datasets is relatively small. Many datasets providing ED $\rho$ are concentrated in the field of material. For instance, MP (Jain et al., 2013) provides approximately 122K PBE-accuracy ED datas, and ECD (Chen et al., 2025) contains 140K PBE-accuracy entries and 7K entries with HSE-accuracy. This ED information provides CHGCAR files. Additionally, for drug-like molecules, QMugs (Isert et al., 2022) and $\nabla^{2}$ (Khrabrov et al., 2024) provide density matrices for 665K and 1.9M molecules, respectively, but do not directly provide ED. Moreover, three ED datasets generated by VASP were used for the equivariant DeepDFT model (Jørgensen and Bhowmik, 2022), including QM9-VASP, but they have few molecules. As atomic-level MLFFs are fully researched, such methods gradually face bottlenecks, highlighting the increasing need for large-scale molecular datasets that incorporate ED and comprehensive benchmarks to facilitate subsequent research. In this study, we constructed a dataset named EDBench, which contains nearly 3.4 million molecules, including ED (CUBE files) and common quantum chemical properties. As shown in Table 1, compared with other datasets, EDBench provides the largest known ED benchmark dataset, characterized by its large scale and comprehensive content.
@@
Category Ours Classical quantum chemistry databases and extensions Molecular dynamics Pharmaceutical Material
Datasets EDBench QM7-X QM9 QH9 QM9- VASP PubChem QC WS22 MD17 QMugs $\nabla^{2}$ QMMD ECD MP
Source PCQM4Mv2 GDB-13 GDB-17 QM9 QM9 PubChem - - CHEMBL MOSES ICSD Magten ICSD
Molecules 3,359,472 7K 134K 130K 134K 85M 10 8 665K 1.9M 1.2M 140K 577K
Element Count 22 6 5 5 5 12 4 4 11 8 - - -
Calc Method (Basis Set/ XC-Function) B3LYP, 6- 31G**/+G** PBE0+ MBD B3LYP, 6-31G (2df,P) +G4MP2 B3LYP, Def2- SVP PBE B3LYP/6-31G* //PM6 PBE0/ 6-31G* PBE+ vdW-TS GFN2-x TB+$\omega$B97 X-D/def 2-SVP $\Omega$B97X- D/Def2- SVP PAW- PBE PBE/HSE 06,GGA+ U/Monkho rst-Pack PBE,GGA /GGA+U, SCAN/R2 SCAN
Electron density $\rho$checkmark-CUBE $\times$ $\times$ $\times$ $\checkmark$-CHG CAR $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$-CHG CAR $\checkmark$(122K)- CHGCAR
Total Energy $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$
NO Occupation $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Canonical Orbit $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$
Dipole Moment $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$
E-Structure $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Software/Tool Psi4 FHI-aims Gaussian PySCF VASP GAMESS Gaussian ORCA/ FHI-aims Psi4 Psi4 VASP VASP VASP

### v2

@@ Abstract
Existing molecular machine learning force fields (MLFFs) generally focus on the learning of atoms, molecules, and simple quantum chemical properties (such as energy and force), but ignore the importance of electron density (ED) $\rho(r)$ in accurately understanding molecular force fields (MFFs). ED describes the probability of finding electrons at specific locations around atoms or molecules, which uniquely determines all ground state properties (such as energy, molecular structure, etc.) of interactive multi-particle systems according to the Hohenberg-Kohn theorem. However, the calculation of ED relies on the time-consuming first-principles density functional theory (DFT), which leads to the lack of large-scale ED data and limits its application in MLFFs. In this paper, we introduce EDBench, a large-scale, high-quality dataset of ED designed to advance learning-based research at the electronic scale. Built upon the PCQM4Mv2, EDBench provides accurate ED data, covering 3.3 million molecules. To comprehensively evaluate the ability of models to understand and utilize electronic information, we design a suite of ED-centric benchmark tasks spanning prediction, retrieval, and generation. Our evaluation of several state-of-the-art methods demonstrates that learning from EDBench is not only feasible but also achieves high accuracy. Moreover, we show that learning-based methods can efficiently calculate ED with comparable precision while significantly reducing the computational cost relative to traditional DFT calculations. All data and benchmarks from EDBench will be freely available, laying a robust foundation for ED-driven drug discovery and materials science.
@@ 1 Introduction
With the widespread adoption of deep learning in molecular dynamics (MD) simulations, machine learning force fields (MLFFs) have become efficient and promising computational tools, significantly advancing research in physics, chemistry, biology, and materials science (wu2023applications, ; kabylda2023efficient, ; frank2024euclidean, ). State-of-the-art MLFFs methods typically employ geometric deep learning to model atomic interactions within molecules, a strategy that has proven to be effective (zheng2025data, ). These models are generally built upon many-body interactions at the atomic level, including one-body (atomic attributes such as types and coordinates (satorras2021n, )), two-body (interatomic distances (schutt2018schnet, ; kabylda2023efficient, )), three-body (bond angles (gasteiger2020fast, ; schutt2021equivariant, )), four-body (torsions (gasteiger2021gemnet, ; liu2022spherical, ; wang2024enhancing, ) and improper torsions (zheng2025data, )), and five-body interactions (wang2023efficiently, ).
@@
Although existing MLFFs have demonstrated great potential in modeling molecular force fields (MFFs), they primarily focus on capturing coarse-grained, atom-level many-body interactions (liu2023symmetry, ; zheng2025data, ), while often overlooking the critical role of microscopic electron distribution in understanding molecular interactions (schuch2009computational, ; gao2019real, ; pederson2022machine, ). It is well known that the spatial distribution of electrons directly influences the interactions between atoms within a molecule, providing the most direct and fundamental information for interpreting molecular force fields (hapala2016mapping, ). Electron density (ED), as a fundamental physical quantity in quantum mechanics that describes the spatial distribution of electrons, offers a more fine-grained and physically grounded representation of molecular systems according to Hohenberg–Kohn (HK) theorem (PhysRev.136.B864, ). Therefore, explicitly incorporating ED into the modeling process holds promise for bridging the gap between microscopic electronic behavior and macroscopic force fields, further improving both the accuracy and generalization of MLFFs. Therefore, as illustrated in Figure 1(a), the primary objective of this work is to advance current MLFFs beyond the atom-level learning paradigm toward electron-level modeling, enabling a more accurate and physically grounded description of molecular interactions. We provide a detailed description of the background in Appendix A.1 to enhance the reader’s understanding.
@@
However, advancing MLFFs toward an electron-level understanding faces two major challenges: (i) the lack of large-scale, high-quality ED datasets, which are essential for pretraining and could fundamentally reshape the paradigm of MLFFs modeling. (ii) the absence of an ED-centric benchmark to systematically explore the feasibility and effectiveness of ED-based modeling frameworks. Specifically, the acquisition of ED data can be categorized into two approaches: experimental methods (such as X-ray diffraction (casati2016putting, ; kasai2018x, ), electron diffraction (gruene2021establishing, )) and theoretical calculation methods. Due to the reliance on expensive physical equipment, experimental methods inevitably limit data acquisition, making theoretical methods more popular. Theoretical calculations typically use density functional theory (DFT) (bartolotti1996introduction, ; medvedev2017density, ), the most common approach, to compute the ED of molecules. Although DFT does not depend on specialized observation equipment, its calculations are highly computationally intensive and time-consuming, making the acquisition of large-scale, high-quality ED datasets particularly difficult. In addition, the MLFFs community is still in the early stages of learning effective representations from ED, which makes the development of an ED-based evaluation protocol particularly important for the rapid advancement of ED representation learning.
@@
To address the two key challenges outlined above, we introduce EDBench, a large-scale and high-fidelity dataset of ED, as shown in Figure 1(b). Following Jacob’s ladder (sousa2007general, ), as shown in Figure 1(c), we adopt higher-rung hybrid functionals as the underlying DFT methods to ensure the quality of the EDBench dataset. Specifically, we use B3LYP function combined with 6-31G**/+G ** basis set and run over 205,000 core-hours on a high-performance computer to generate EDBench. EDBench comprises 3,359,472 drug-like molecules with corresponding ED distributions and a comprehensive set of quantum chemical properties, including energy components, orbital energies, and multipole moments, thus providing a solid foundation for systematically investigating the role of ED in molecular modeling. To enable rigorous benchmarking and model development, we further design an ED-centric benchmark suite covering three task categories: (i) quantum property prediction, including four core tasks—energy components prediction (ED5-EC), orbital energy estimation (ED5-OE), multipole moment regression (ED5-MM), and open-/closed-shell classification (ED5-OCS)—to evaluate how well ED alone can serve as a sufficient descriptor for inferring fundamental quantum properties; (ii) cross-modal retrieval between molecular structures and ED (ED5-MER), designed to probe the mutual consistency and representational alignment between structural and density spaces, which is critical for density-based force field construction and virtual screening; and (iii) ED prediction from molecular structures (ED5-EDP), aimed at approximating DFT-level density fields at significantly reduced computational cost, thereby enabling scalable quantum-aware modeling. Finally, we evaluate several state-of-the-art deep learning models on the proposed benchmark, offering the first large-scale assessment of ED understanding in data-driven systems.
@@ 2.1 Density Functional Theory (DFT)
The quantum mechanical description of many-electron systems is one of the core issues in modern physics and chemistry. Schrödinger equation (Schr1926An, ) as the fundamental equation of quantum mechanics, is challenging to solve directly. Consequently, researchers introduced various wave function-based approximation methods to simplify the problem, such as, Born–Oppenheimer (Combes1981, ) and Hartree-Fock method (szaba1996, ). Those methods scale with the number of electrons $n$ as $\mathcal{O}(n^{4})$ or more, its computational cost remains prohibitive for large polyatomic molecules (More details about the development of quantum mechanics see Appendix A.2.1). In contrast, Density Functional Theory (DFT) is more suitable for complex systems due to its lower computational cost ($\mathcal{O}(n^{3})$) and incorporation of electron correlation effects (gaitan2009, ). The core concept of DFT is to use electron density (ED) as the fundamental variable instead of the wave function. The Hohenberg-Kohn theorem is the cornerstone of DFT (PhysRev.136.B864, ), which states that the external potential field and the ground-state energy can be completely determined by ED. Thus, by solving for the ED distribution $\rho(r)$ that achieves the lowest energy, the properties of the stable system can be confirmed. The ED $\rho(r)$ can be expressed as:
@@
where $\rho_{\alpha}(\mathbf{r})$ and $\rho_{\beta}(\mathbf{r})$ are the density of $\alpha$-spin electrons and $\beta$-spin electrons. This concept is concretely realized in the Kohn-Sham equations, which transforms the polyelectron system with interactions into single-electron system without interaction, and adds interactions among electrons to exchange-correlation potential (KohnSham1965, ). The Kohn-Sham equations is shown as:
@@
The solution of the Kohn-Sham equations is typically achieved through self-consistent field (SCF) iterations, as shown in the figure 1(b). Initially, a set of initial electron densities $\rho(\mathbf{r})$ is selected, and the effective potential $V_{\text{eff}}(\mathbf{r})$ is calculated based on this initial guess. The Kohn-Sham equations are then solved to obtain new single-electron orbital wave functions $\psi_{i}(\mathbf{r})$ and energies $\varepsilon_{i}$, which are used to update the ED $\rho(\mathbf{r})$. This process is repeated until convergence is achieved, yielding the ED $\rho(\mathbf{r})$ and simultaneously stabilizing the total energy $E$. The choice of the exchange-correlation functional determines the accuracy of DFT, as shown in Figure 1(c). Common approximations include the Local Density Approximation (LDA) (PhysRevLett.77.3865, ), the Generalized Gradient Approximation (GGA) (Becke1988, ), and hybrid functionals (B3LYP) (PhysRevB.37.785, ). In this paper, we used B3LYP functional combined with 6-31G**/+G** basis set, which achieves a balance between precision and efficiency(Details on basis set and functional see Appendix A.2.3).
@@ 2.2 Molecular Geometry Learning in Quantum Chemistry
Geometric Deep Learning (GDL) has become a dominant approach for modeling machine learning force fields (MLFFs), primarily focusing on atom-level information such as atomic attributes and interatomic interactions. Specifically, GDL models are built upon first-order atomic features, including atom types and 3D coordinates (satorras2021n, ; cui2024geometry, ). To capture geometric relationships while preserving physical consistency, GDL methods incorporate symmetries such as rotational and translational invariance in 3D space (musil2021physics, ; gebauer2022inverse, ). Consequently, a wide range of models have been developed with built-in invariance or equivariance to Euclidean group E(3) (satorras2021n, ) or special Euclidean group SE(3) (ganea2021geomol, ; schneuing2024structure, ), ensuring that predictions are physically meaningful. Given that atomic interactions—such as bonding, repulsion, and van der Waals forces—play a crucial role in molecular fields, modern GDL methods further incorporate second-order geometric features, including interatomic distances (unke2019physnet, ; fuchs2020se, ), bond types (fang2022geometry, ), and spatial neighborhood structures (zhang2024geometry, ). To more precisely capture local structural features, some approaches even extend to higher-order geometric relations such as bond angles (three-body interactions) (gasteiger2020fast, ; schutt2021equivariant, ) and torsional angles (four-body interactions) (gasteiger2021gemnet, ; liu2022spherical, ; wang2024enhancing, ), thereby improving the expressiveness and accuracy of force field modeling. In contrast to prior works that focus primarily on atom-level representations, our proposed EDBench introduces a large-scale dataset of electronic density (ED), laying the foundation for extending molecular modeling from the atomic scale to the electronic scale. It also provides a new platform and evaluation benchmark for developing GDL methods tailored to electronic structure modeling.
@@ 2.3 Quantum Chemistry (QC) Datasets
To more efficiently predict quantum chemical properties and enhance applications using machine learning models, numerous QC datasets have been constructed to provide rich and reliable data support, as shown in Table 1. QM7 (rupp, ) and QM9 (ramakrishnan2014quantum, ) are classic QC datasets, built based on GDB-13 (blum, ) and GDB-17 (Ruddigkeit2012, ), respectively. QM7-X (Hoja2021, ) covers up to 42 physicochemical properties. PubQChemQC (doi:10.1021/acs.jcim.3c00899, ) offers 85 million ground-state molecular structures and HOMO-LUMO gaps. Additionally, MD17 (doi:10.1126/sciadv.1603015, ), MD22 (chmiela2023accurate, ) and WS22 (Pinheiro2023, ) offer force data from molecular dynamics trajectories. Based on QM9 dataset, QH9 (Yu2023, ) focuses on the Hamiltonian matrix, and MultiXC-QM9 (Nandi2023, ) includes information on chemical reactions. Compared with QC datasets that provide energy (Jie2019, ) and force, the proportion of existing ED datasets is relatively small. Many datasets providing ED $\rho$ are concentrated in the field of material. For instance, MP (10.1063/1.4812323, ) provides approximately 122K PBE-accuracy ED datas, and ECD (chen2025ecd, ) contains 140K PBE-accuracy entries and 7K entries with HSE-accuracy. This ED information provides CHGCAR files. Additionally, for drug-like molecules, QMugs (Isert2022, ) and $\nabla^{2}$DFT (2024arXiv240614347K, ) provide density matrices for 665K and 1.9M molecules, respectively, but do not directly provide ED. Moreover, three ED datasets generated by VASP were used for the equivariant DeepDFT model (Jorgensen2022, ), including QM9-VASP, but they have few molecules. As atomic-level MLFFs are fully researched, such methods gradually face bottlenecks, highlighting the increasing need for large-scale molecular datasets that incorporate ED and comprehensive benchmarks to facilitate subsequent research. In this study, we constructed a dataset named EDBench, which contains nearly 3.4 million molecules, including ED (CUBE files) and common quantum chemical properties. As shown in Table 1, compared with other datasets, EDBench provides the largest known ED benchmark dataset, characterized by its large scale and comprehensive content. We also summarize physical and chemical properties of different quantum datasets in Appendix A.3.
@@
[…]